After the ~36 nmi (~3.75 hr) transit to Site 1115 (ACE-15A), a positioning beacon was deployed at the Site 1115 GPS coordinates at 1145 hr on 16 July.

SITE 1115

Hole 1115A

After preparing an APC coring assembly and lowering it to the seafloor, we spudded Hole 1115A at 1510 hr on 16 July. The single APC core taken at Hole 1115A recovered 4.43 m; therefore, the seafloor was 1149.6 mbsl. The bit was pulled clear of the mudline, immediately ending Hole 1115A, Core 1H was recovered, and the pipe was spaced out for spudding Hole 1115B.

 Hole 1115B

At 1540 hr, Hole 1115B was spudded and Core 1H recovered 7.14 m, establishing the seafloor depth at 1148.8 mbsl. Cores 1H through 23H were taken to 216.2 mbsf. Core orientation (tensor tool) was performed on Cores 4H–23H, and Adara temperature measurements were taken after Cores 3H (26.2 mbsf), 5H (45.2 mbsf), 7H (64.2 mbsf), and 9H (83.2 mbsf) were recovered. APC coring was suspended after Core 23H when 70,000 lb overpull was required to extract the core barrel from the formation. APC core recovery was 225.67 m (104%).

Use of the XCB began with Core 24X and continued through Core 31X to a depth of 1453.4 m or 293.1 mbsf. A single DVTP temperature measurement was taken after Core 23X at a depth of 225.7 mbsf. Hole conditions remained excellent as did the ROP, which averaged 19.2 and 38.4 m/hr respectively on the last two cores. Recovery for the XCB portion of the hole was 80%. The hole was terminated after Core 31X because the formation had become indurated enough for successful RCB coring and because this section would have to be drilled again later in Hole 1115C. Total recovery for the hole was 286.84 m (98%).

The drill string was pulled clear of the seafloor at 1415 hr on 17 July, and the bit cleared the rig floor at 1615 hr that same day, ending Hole 1115B.

 Hole 1115C

While we assembled the RCB BHA, the ship was offset 30 m to the west. The drill string was lowered to the seafloor and Hole 1115C was spudded at 2130 hr on 17 July. A seafloor depth of 1148.7 mbsl was established, based on a reduction in drill string weight. With a center bit in place, the hole was drilled without coring to 283.2 mbsf. The hole was flushed clean and the center bit recovered, preparing the way for RCB coring.

The RCB coring began at 0400 hr on 18 July 1998 and continued through Core 54R to 802.5 mbsf. Coring was terminated when the scientists considered that the primary scientific objectives had been achieved. Two zones of low recovery were encountered: 7% from 523.0 to 542.2 mbsf and from 1% to 2% from 628.8 to 657.8 mbsf. Bentonite gel mud sweeps (30 bbl each) were pumped every third core, and hole conditions remained excellent throughout the coring operations. Penetration rates were quite variable, ranging from 5.8 m/hr to 57.6 m/hr and averaged 18.0 m/hr. Total RCB core recovery was 291.63 m (56.2%).

A wiper trip up to 99.7 mbsf and then back to the bottom identified tight spots at 701.7 and 762.7 mbsf where the driller observed 40,000 lb overpull. The bit encountered 15 m of fill in the bottom of the hole, which we decided not to spend the time washing, reaming, and circulating out of the hole. The bit was released, the hole displaced with 215 bbl of sepiolite mud, and the drill pipe positioned at 99.7 mbsf in preparation for logging.

We began rigging up for logging at 0500 hr on 21 July 1998, and the first logging run consisted of the triple combo. The suite of tools reached a depth of 783.7 mbsf, which was 4.0 m above where the bit was released. The second tool string (FMS-sonic) also reached the same depth. Two logging passes were made with the FMS before pulling it out of the hole. The third logging string run was the ultrasonic borehole imager (UBI). This tool was put aboard by Schlumberger in lieu of a second borehole televiewer (BHTV) tool. The UBI is another form of borehole imager used extensively in industry, but this was the first time this tool had been deployed for ODP. The tool reached the same depth as the previous tool runs but, unfortunately, did not function correctly. It was recovered without obtaining any data. The final logging run was with the WST, but it failed to pass a bridge at 564.7 mbsf (probably because of its light weight). An air gun was deployed using the aft port side aft crane, and several stations were successfully occupied until the logging run had to be aborted because of an expected helicopter landing. During helicopter landings, the cranes must not be deployed for safety reasons. At 0600 hr, the WST was pulled up inside the pipe, the air gun was rigged down, and the crane used to deploy the air gun was retracted. The helicopter was bringing out a reporter from the Australian Broadcasting Commission (ABC) and a roustabout who had recovered from illness after having been evacuated from the ship earlier in the leg. We were notified by the ship’s agent in PNG that the helicopter would depart Alotau at ~0630 hr on 22 July and would arrive on location within 30 min. By 0730 hr the helicopter had still not arrived and the radio officer made several unsuccessful attempts to contact the helicopter via radio. He was able to hear conversations between other airborne helicopters but was not able to contact the helicopter destined for the JOIDES Resolution. At ~0845 hr the ship received a phone call from the agent that the helicopter had gone to the wrong site coordinates. Unable to locate the ship the helicopter had returned to Alotau. At ~0900 hr we resumed VSP logging operations. Subsequent VSP stations were not as successful as those earlier because of an unidentified source of noise in the system. After attempts to isolate and fix the cause of the noise, the VSP was terminated. The WST was pulled out of the hole and the air guns rigged down. All logging operations were completed, and the tools rigged down by 1345 hr on 22 July 1998.

The open-ended drill pipe was lowered to 483.7 mbsf, and the hole was displaced with 100 bbl of 10.5-ppg weighted gel mud. The drill string was pulled clear of the seafloor at 1630 hr, and the positioning beacon was recovered at 1745 hr. After recovering the drill string, the BHA drill collars and associated subs were subjected to the routine end of leg inspection for cracks (none were identified). Once the BHA inspection was completed, the final thruster was raised and we began the transit to Site 1116 (proposed site ACE-16A) at 2230 hr on 22 July 1998.


To 180 Site 1116

To 180 Table of Contents